# Neural network embeddings recover value dimensions from psychometric survey items on par with human data

### v1

@@ Abstract
This study introduces "Survey and Questionnaire Item Embeddings Differentials" (SQuID), a novel methodological approach that enables neural network embeddings to effectively recover latent dimensions from psychometric survey items. We demonstrate that embeddings derived from large language models, when processed with SQuID, can recover the structure of human values obtained from human rater judgments on the Revised Portrait Value Questionnaire (PVQ-RR). Our experimental validation compares multiple embedding models across a number of evaluation metrics. Unlike previous approaches, SQuID successfully addresses the challenge of obtaining negative correlations between dimensions without requiring domain-specific fine-tuning. Quantitative analysis reveals that our embedding-based approach explains 55% of variance in dimension-dimension similarities compared to human data. Multidimensional scaling configurations from both types of data show fair factor congruence coefficients and largely follow the underlying theory. These results demonstrate that semantic embeddings can effectively replicate psychometric structures previously established through extensive human surveys. The approach offers substantial advantages in cost, scalability and flexibility while maintaining comparable quality to traditional methods. Our findings have significant implications for psychometrics and social science research, providing a complementary methodology that could expand the scope of human behavior and experience represented in measurement tools.
@@
Max Pellert1, Clemens M. Lechner2, Indira Sen3, Markus Strohmaier2,3,4 1Barcelona Supercomputing Center 2GESIS – Leibniz Institute for the Social Sciences 3University of Mannheim 4Complexity Science Hub Vienna Correspondence: max.pellert@bsc.es
@@ 1 Introduction
Psychometrics, the field of study in psychology that is crucially concerned with measurement of human attributes such as skills, personality, or values, is undergoing a revolution with the integration of novel NLP technologies to complement and extend traditional questionnaire-based workflows. Proposed ideas range from the assessment of large language models via "AI Psychometrics" Pellert et al. (2024) to "Machine Psychology" Hagendorff (2023) or, more conservatively, to use LLMs as assisting tools in human psychometrics to aid comparably cheap, in-silico questionnaire item creation Götz et al. (2023) or automated coding Nilsson et al. (2025). For the closely related task of synthetic data generation Hommel et al. (2022), LLMs have been used to simulate segments of the population to match survey results Argyle et al. (2023) over time Ahnert et al. (2025) and to replicate experimental studies Aher et al. (2023); Dillion et al. (2023).
@@
Although the potential is large, critical evaluation has shown various limitations of existing attempts to marry NLP technologies with social scientific applications such as psychometric measurement Löhn et al. (2024); Sühr et al. (2024). We review selected representative approaches by other researchers and discuss the problems they face. Then, we present our novel methodological advancement of "Survey and Questionnaire Item Embeddings Differentials" (SQuID) and demonstrate with the Revised Portrait Value Questionnaire (PVQ-RR) how it allow us to overcome issues that we identified.
@@ 2 Related Work
A number of recent studies have proposed using embeddings of questionnaire items from psychometric inventories—or various aggregates of items. These embedding vectors can then be used to create similarity matrices consisting of pairwise similarities between all items or item aggregates. These, in turn, can be analyzed with factor-analytic or similar dimension-reduction methods in much the same way that researchers have analyzed correlation matrices resulting from human subjects’ responses to psychometric inventories. On a fundamental level that also ensures full reproducibility, Cutler and Condon (2023) successfully replicated a classical human subject study on personality traits using an embedding approach. The authors created templates: "Those close to me say I have a [MASK][MASK] and [TERM] personality." [TERM] is inserted from a list of personality descriptors containing entries such as "condescending", the constructed statement is passed through the model and the token embeddings of the two [MASK] tokens are extracted and averaged. The results of principal component analysis on that averaged embedding were reasonably similar to results of a classical study on human ratings of personality descriptors Saucier and Goldberg (1996) with 3 out of 5 factors adequately well recovered. The authors used a less systematic and more combinatorial approach of running the same experiment with a large number of different models, all versions of BERT, to find the best-performing one according to factor congruence scores with the original human rater judgments. Interestingly, a model finetuned for Natural Language Inference (NLI) tasks, the head of which was in fact discarded for this setup (as only the two masked token embeddings are extracted), performed best. Cutler and Condon did not provide a systematic investigation on why this specific model worked best. Moreover, because of the structure of the templates containing personality descriptors, their approach is largely limited to personality inventories for which this structure is adequate. While their replication study did not provide an explicit acknowledgment of the generality of the approach, it presented pioneering evidence of the power of neural embeddings in emulating human rater judgments.
@@
Researchers have also taken a more macro-perspective and worked with embeddings of data sets that collect a number of different personality questionnaires. Wulff and Mata (2025) finetune a BERT based model on 200,000 pairs of items and their empirical correlations to create a domain-specific embedding model. Laudably, they share their model publicly on the Hugging Face Model Hub  [^1] https://huggingface.co/dwulff/mpnet-personality and we include it in our analyses. They use their model to embed a large number of different personality items, scales and construct names. These embeddings allow them to predict empirical similarities between items reasonably well, to detect overlaps between scales and to suggest more parsimonious taxonomies. Relying on a very similar base model, Hommel and Arslan (2025) used finetuning to create an embedding model that allows them to establish inter-item correlations, scale reliability and inter-scale correlations to some degree. Unfortunately, their model is not openly shared and can therefore not be used by us or by other researchers.
@@
These approaches are generalized by Feraco and Toffalini (2025) by building on the before-mentioned "mpnet-personality" model to create a pipeline module that allows researchers to run in-silico validity checks on item candidates to identify problems in a cheap and easy way before any empirical data from human raters is collected. Similar, "pseudo-factor analysis" was recently proposed as a pre-testing approach that allows evaluating the semantic structure of personality inventories (and potentially other psychological constructs) Guenole et al. (2025). Like the approach by Feraco and colleagues, pseudo-factor analyis is based on embedding items or parcels of items and computing similarity matrices between embedding vectors, which can then be factor-analysed in the same way as inter-item correlations of personality ratings from human respondents are traditionally factor analyzed. Guenole and colleagues illustrated their approach with two widely used personality inventories, reporting that pseudo-factor analyses was able to replicate empirical findings on the structure of these inventories obtained on human subject reasonably well, at least when using (the rather permissive) oblique rotation. Despite these recent advances, it remains unclear which of the many different models and strategies for creating and combining embeddings on different items or item parcels works best, let alone why.
@@
Different to these prior attempts, we aim for a general methodology that can readily be applied post-hoc to embeddings of diverse contexts also different to personality derived from any model without re-annotating training corpora and finetuning. As described in Section 3.5, we develop "Survey and Questionnaire Item Embeddings Differentials" (SQuID) as a general method that fulfills these constraints and allows to obtain negative correlations.
3 Experimental Setup
@@ 3.1 Fundamental assumptions when adapting psychometric workflows to LLMs
3.2 The theory of values
All the works discussed in Section 2 have so far been in the domain of questionnaires for personality traits and attendant inventories (e.g., Big Five and HEXACO), which are a natural starting point because of the long research tradition of the ’lexical approach’ investigating the semantic structure of personality-descriptive adjectives and statements Guenole et al. (2025). However, for embedding-based approaches to be a truly valuable addition to the psychometric toolkit, their viability needs to be demonstrated on additional constructs beyond personality traits. In our study, we therefore aim to go beyond personality inventories and choose a widely used and well-established survey for the important domain of values. According to Schwartz’s theory of basic human values, values are trans-situational goals that serve as guiding principles in people’s lives, expressing distinct motivational concerns. The "Revised Portrait Value Questionnaire" (PVQ-RR) Schwartz et al. (2012) contains 57 items such as "It is important to him/her to form his/her views independently.", "It is important to him/her that his/her country is secure and stable.", "It is important to him/her to have a good time.", which can be aggregated to 19 fine-grained value dimensions such as "Self-direction Thought", "Security-Societal" and "Hedonism" (3 items per dimension). Human respondents rate how similar they view themselves to the person described in the item on a 6-point scale ranging from "not like me at all" to "very much like me". The survey scoring rules that specify the dimension that each item belongs to and can be aggregated to through averaging have been developed in a theory-guided way that has been cross-culturally validated and refined since the 1980ies with human rater judgments from 49 countries Schwartz and Cieciuch (2022).
@@ 3.3 Multidimensional Scaling and the Centrality of Negative Correlations
Importantly, negative correlations between items play a very important role in Schwartz’s theory of the content and structure of human values. Schwartz’s values are structured in a circumplex model based on the compatibility and conflict between underlying motivations, whereby values that motivational oppose each other are placed on opposite positions of the circle and correlate negatively with each other (e.g. Conformity vs. Self-direction), whereas closely related values that are motivational compatible with each other are positively correlated and located close to each other on the same side of the circle (e.g. Achievement and Power). The individual values can be further aggregated to four higher-order dimensions corresponding to four segments of the circle that emerge with great regularity in data from human respondents Lechner et al. (2024).
This circular structure is uncovered using a dimension-reduction technique that is different from principal component analysis and factor analysis as typically used for personality inventories: Multi-dimensional scaling (MDS) Jung and Takane (2015); Groenen and Borg (2015). Schwartz’s theory and attendant inventories were historically developed in close relationship with this technique. MDS maps high-dimensional objects such as vectors of human rater judgments or semantic embeddings to points in a lower-dimensional, typically 2D, space. The technique works in such a way that points corresponding to similar objects are located close together, while dissimilar objects are located far apart, thereby preserving distances.
The MDS configurations that are computed from human rater judgments confirm those relationships between values as postulated by the theory. We can therefore empirically validate the theoretically assumed circular placement of values using MDS with human rater judgments Schwartz and Cieciuch (2022), while, until now, it was an open question if the same is possible with an embedding approach. The paramount importance of negative correlations between items renders Schwartz’s value theory and the PVQ-RR an ideal (though challenging) case to refine embeddings-based approaches such that they reproduce negative correlations between items foundational to value theory.
@@ 3.4 Embedding Models
The speed of development makes it impossible to provide a systematic model review that stays state-of-art long enough to even account for the length of typical publication cycles. For this reason, we instead employ strategic selection to include models in our analyses that are representative for specific model classes and types. This way, we can provide some guidance on which models typically to look for in this application area also in the future.
First, we choose the currently best performing open model for semantic textual similarity (STS) tasks in "MTEB: Massive Text Embedding Benchmark" Muennighoff et al. (2023), Linq-Embed-Mistral (7.11B, prompt, Choi et al. (2024)). In addition, we choose the currently best performing closed API model in the benchmark, gemini-embedding-exp-03-07 (parameter count not disclosed, no prompt). Furthermore, we include two smaller models created with innovative training and data strategies that allow them to perform well for their size, jina-embeddings-v3 (572M, no prompt, Sturua et al. (2024)) and KaLM-embedding-multilingual-mini-instruct-v1.5 (494M, prompt, Hu et al. (2025)). Finally, we include the model mpnet-personality (109M, no prompt) that, as discussed before, was created through domain specific finetuning. Our approach does not need extensive prompt engineering, any sophisticated token probability extraction strategies or accounting for stochasticity of outputs. In a growing social scientific literature of using LLMs for empirical investigations, instead of interacting with models on the textual output level we show the advantages of working on the fundamental level of semantic embeddings. We use the exact wording from the PVQ-RR questionnaire specifications (see Subsection A.7) for those models that allow to include a prompt [^2] We run all models on CPU to embed all 57 short texts of the items in less than 20 minutes on a ThinkPad T480.. The PVQ-RR is available in gendered version using different pronouns. We account for this feature by creating an average of the female and male item embeddings in order to reduce possible gender biases.
@@ 3.5 Survey and Questionnaire Item Embeddings Differentials (SQuID)
Working with raw embeddings as directly retrieved from the models, we faced the same problem of missing negative correlations as the other researchers whose work we discussed in Section 2 (see the similarity matrix on the left of Figure 4). As we will use MDS for dimensionality reduction which builds on preserving distances, it was especially important for us to find a way to make the signal of negative correlations stronger.
As discussed before, the root of the problem lies in the high-dimensionality of the item embeddings that encode many common linguistic features of natural language use which artificially boost similarity between items. To create a vector that captures this unneeded information, we take an average embedding over all items (a "questionnaire embedding") and subtract it from each of the item embeddings: $\mathbf{y}_{i}-\overline{\mathbf{y}}$ for item $i$, where $\overline{\mathbf{y}}$ is an average over all 57 items of the PVQ-RR: $\overline{\mathbf{y}}=\frac{1}{57}\sum_{i=1}^{57}\mathbf{y}_{i}$. This straightforward method is inspired by the the early examples of vector arithmetic demonstrated with Word2Vec Mikolov et al. (2013) and the more recent idea of community embeddings Waller and Anderson (2021).
As can be seen in the right hand part of Figure 4, this approach is simple yet very effective in making the signal of negative correlations appear without any domain-specific finetuning or other adaption procedures. SQuID works with any model and embeddings of arbitrary dimensions, offering a high degree flexibility. We believe that a substantial part of the under-performance of embedding approaches compared to human data in other studies, for example in Wang et al. (2024), is due to methodological reasons and can be overcome with SQuID.
@@
Table 1: Internal consistency as measured through Cronbach’s alpha for each of the 19 value dimensions and the average over dimensions. Rowname abbreviations refer to the models covered in Subsection 3.4 (all embeddings were SQuID treated). We report a random baseline of embeddings created from a normal distribution with mean zero and unit variance (averages over 10 000 samples). Human data is retrieved from Table S2 in SI of Schwartz and Cieciuch (2022). Linq-Embed-Mistral is generally the best performing embedding model with an average value for Cronbach’s alpha that is above even the human benchmark. Values close to 0.8 or above for most value dimension point to adequate internal consistency (with notable exceptions such as HUM/humility, which is also low in the human data). For fully spelled out dimension names in the column header see Figure 4.

### v2

@@ Abstract
We demonstrate that embeddings derived from large language models, when processed with "Survey and Questionnaire Item Embeddings Differentials" (SQuID), can recover the structure of human values obtained from human rater judgments on the Revised Portrait Value Questionnaire (PVQ-RR). We compare multiple embedding models across a number of evaluation metrics including internal consistency, dimension correlations and multidimensional scaling configurations. Unlike previous approaches, SQuID addresses the challenge of obtaining negative correlations between dimensions without requiring domain-specific fine-tuning or training data re-annotation. Quantitative analysis reveals that our embedding-based approach explains 55% of variance in dimension-dimension similarities compared to human data. Multidimensional scaling configurations show alignment with pooled human data from 49 different countries. Generalizability tests across three personality inventories (IPIP, BFI-2, HEXACO) demonstrate that SQuID consistently increases correlation ranges, suggesting applicability beyond value theory. These results show that semantic embeddings can effectively replicate psychometric structures previously established through extensive human surveys. The approach offers substantial advantages in cost, scalability and flexibility while maintaining comparable quality to traditional methods. Our findings have significant implications for psychometrics and social science research, providing a complementary methodology that could expand the scope of human behavior and experience represented in measurement tools.
@@
Max Pellert1, Clemens M. Lechner2, Indira Sen3, Markus Strohmaier3,4 1Barcelona Supercomputing Center 2Jam Technologies GmbH 3University of Mannheim 4GESIS - Leibniz Institute for the Social Sciences Correspondence: max.pellert@bsc.es
@@ 1 Introduction
Psychometrics, the field of study in psychology that is crucially concerned with measurement of human attributes such as skills, personality or values, is undergoing a revolution with the integration of novel NLP technologies to complement and extend traditional questionnaire-based workflows. Proposed ideas range from the assessment of large language models via "AI Psychometrics" Pellert et al. (2024) to "Machine Psychology" Hagendorff (2023) or, more conservatively, to use LLMs as assisting tools in human psychometrics to aid comparably cheap, in-silico questionnaire item creation Götz et al. (2023) or automated coding Nilsson et al. (2025). For the closely related task of synthetic data generation Hommel et al. (2022), LLMs have been used to simulate segments of the population to match survey results Argyle et al. (2023) over time Ahnert et al. (2025) and to replicate experimental studies Aher et al. (2023); Dillion et al. (2023).
@@
Although the potential is large, critical evaluation has shown various limitations of existing attempts to marry NLP technologies with social scientific applications such as psychometric measurement Löhn et al. (2024); Sühr et al. (2024). We review selected representative approaches by other researchers and discuss the problems they face. Then, we present our novel methodological advancement of "Survey and Questionnaire Item Embeddings Differentials" (SQuID) and demonstrate with the Revised Portrait Value Questionnaire (PVQ-RR) how it allows us to overcome issues that we identified.
@@ 2 Related Work
A number of recent studies have proposed using embeddings of questionnaire items from psychometric inventories—or various aggregates of items. These embedding vectors can then be used to create similarity matrices consisting of pairwise similarities between all items or item aggregates. These, in turn, can be analyzed with factor-analytic or similar dimension-reduction methods in much the same way that researchers have analyzed correlation matrices resulting from human subjects’ responses to psychometric inventories. On a fundamental level that also ensures full reproducibility, Cutler and Condon (2023) successfully replicated a classical human subject study on personality traits using an embedding approach. The authors created templates: "Those close to me say I have a [MASK][MASK] and [TERM] personality." [TERM] is inserted from a list of personality descriptors containing entries such as "condescending", the constructed statement is passed through the model and the token embeddings of the two [MASK] tokens are extracted and averaged. The results of principal component analysis on that averaged embedding were reasonably similar to results of a classical study on human ratings of personality descriptors Saucier and Goldberg (1996) with 3 out of 5 factors adequately well recovered. The authors used a less systematic and more combinatorial approach of running the same experiment with a large number of different models, all versions of BERT, to find the best-performing one according to factor congruence scores with the original human rater judgments. Interestingly, a model finetuned for Natural Language Inference (NLI) tasks, the head of which was in fact discarded for this setup (as only the two masked token embeddings are extracted), performed best. Cutler and Condon did not investigate why this model worked best. Moreover, their template-based approach is largely limited to personality inventories, for which this structure is adequate. While their replication study did not provide an explicit acknowledgment of the generality of the approach, it presented pioneering evidence of the power of neural embeddings in emulating human rater judgments.
@@
Researchers have also taken a more macro-perspective and worked with embeddings of data sets that collect a number of different personality questionnaires. Wulff and Mata (2025) finetune a BERT based model on 200,000 pairs of items and their empirical correlations to create a domain-specific embedding model. Laudably, they share their model publicly on the Hugging Face Model Hub  [^1] https://huggingface.co/dwulff/mpnet-personality and we include it in our analyses. They use their model to embed a large number of different personality items, scales and construct names. These embeddings allow them to predict empirical similarities between items reasonably well, to detect overlaps between scales and to suggest more parsimonious taxonomies. Relying on a very similar base model, Hommel and Arslan (2025) used finetuning to create an embedding model that allows them to establish inter-item correlations, scale reliability and inter-scale correlations to some degree. Unfortunately, their model is not openly shared and therefore cannot be used by us or by other researchers.
@@
These approaches are generalized by Feraco and Toffalini (2025) by building on the before-mentioned "mpnet-personality" model to create a pipeline module that allows researchers to run in-silico validity checks on item candidates to identify problems in a cheap and easy way before any empirical data from human raters is collected. Similarly, "pseudo-factor analysis" was recently proposed as a pre-testing approach that allows evaluating the semantic structure of personality inventories (and potentially other psychological constructs) Guenole et al. (2025). Like the approach by Feraco and colleagues, pseudo-factor analyis is based on embedding items or parcels of items and computing similarity matrices between embedding vectors, which can then be factor-analysed in the same way as inter-item correlations of personality ratings from human respondents are traditionally factor analyzed. Guenole and colleagues illustrated their approach with two widely used personality inventories, reporting that pseudo-factor analyses was able to replicate empirical findings on the structure of these inventories obtained on human subject data reasonably well, at least when using (the rather permissive) oblique rotation. Despite these recent advances, it remains unclear which of the many different models and strategies for creating and combining embeddings on different items or item parcels works best, let alone why.
@@
Different to these prior attempts, we aim for a general methodology that can readily be applied post-hoc to embeddings of diverse contexts also different from personality derived from any model without re-annotating training corpora and finetuning. As described in Section 4.2, we develop "Survey and Questionnaire Item Embeddings Differentials" (SQuID) as a general method that fulfills these constraints and allows obtaining negative correlations.
3 Background
3.1 The theory of values
All the works discussed in Section 2 have so far been in the domain of questionnaires for personality traits and attendant inventories (e.g., Big Five and HEXACO), which are a natural starting point because of the long research tradition of the ’lexical approach’ investigating the semantic structure of personality-descriptive adjectives and statements Guenole et al. (2025). However, for embedding-based approaches to be a truly valuable addition to the psychometric toolkit, their viability needs to be demonstrated on additional constructs beyond personality traits. In our study, we therefore aim to go beyond personality inventories and choose a widely used and well-established survey for the important domain of values. According to Schwartz’s theory of basic human values, values are trans-situational goals that serve as guiding principles in people’s lives, expressing distinct motivational concerns. The "Revised Portrait Value Questionnaire" (PVQ-RR) Schwartz et al. (2012) contains 57 items such as "It is important to him/her to form his/her views independently.", "It is important to him/her that his/her country is secure and stable.", "It is important to him/her to have a good time.", which can be aggregated to 19 fine-grained value dimensions such as "Self-direction Thought", "Security-Societal" and "Hedonism" (3 items per dimension). Human respondents rate how similar they view themselves to the person described in the item on a 6-point scale ranging from "not like me at all" to "very much like me". The survey scoring rules that specify the dimension that each item belongs to and can be aggregated to through averaging have been developed in a theory-guided way that has been cross-culturally validated and refined since the 1980s with human rater judgments from 49 countries Schwartz and Cieciuch (2022).
@@ 3.3 Multidimensional Scaling and the Centrality of Negative Correlations
Importantly, negative correlations between items play a very important role in Schwartz’s theory of the content and structure of human values. Schwartz’s values are structured in a circumplex model based on the compatibility and conflict between underlying motivations, whereby values that motivationally oppose each other are placed on opposite positions of the circle and correlate negatively with each other (e.g. Conformity vs. Self-direction), whereas closely related values that are motivationally compatible with each other are positively correlated and located close to each other on the same side of the circle (e.g. Achievement and Power). The individual values can be further aggregated to four higher-order dimensions corresponding to four segments of the circle that emerge with great regularity in data from human respondents Lechner et al. (2024).
This circular structure is uncovered using a dimension-reduction technique that is different from principal component analysis and factor analysis as typically used for personality inventories: Multi-dimensional scaling (MDS) Jung and Takane (2015); Groenen and Borg (2015). Schwartz’s theory and attendant inventories were historically developed in close relationship with this technique. MDS maps high-dimensional objects such as vectors of human rater judgments or semantic embeddings to points in a lower-dimensional, typically 2D, space. The technique works in such a way that points corresponding to similar objects are located close together, while dissimilar objects are located far apart, thereby preserving distances. The MDS configurations that are computed from human rater judgments confirm those relationships between values as postulated by the theory. We can therefore empirically validate the theoretically assumed circular placement of values using MDS with human rater judgments Schwartz and Cieciuch (2022), while, until now, it was an open question if the same is possible with an embedding approach. The paramount importance of negative correlations between items renders Schwartz’s value theory and the PVQ-RR an ideal (though challenging) case to refine embeddings-based approaches such that they reproduce negative correlations between items foundational to value theory.
4 Experimental Setup
@@ 4.1 Embedding Models
Rapid development makes systematic model reviews quickly outdated. We therefore strategically selected models representative of specific model classes and types. This way, we can provide some guidance on which models typically to look for in this application area also in the future.
First, we choose the currently best performing open model for semantic textual similarity (STS) tasks in "MTEB: Massive Text Embedding Benchmark" Muennighoff et al. (2023), Linq-Embed-Mistral (7.11B, prompt, Choi et al. 2024). In addition, we choose the currently best performing closed API model in the benchmark, gemini-embedding-exp-03-07 (parameter count not disclosed, no prompt). Furthermore, we include two smaller models created with innovative training and data strategies that allow them to perform well for their size, jina-embeddings-v3 (572M, no prompt, Sturua et al. 2024) and KaLM-embedding-multilingual-mini-instruct-v1.5 (494M, prompt, Hu et al. 2025). Finally, we include the model mpnet-personality (109M, no prompt) that, as discussed before, was created through domain specific finetuning. Our approach does not need extensive prompt engineering, any sophisticated token probability extraction strategies or accounting for stochasticity of outputs. In a growing social scientific literature of using LLMs for empirical investigations, instead of interacting with models on the textual output level we show the advantages of working on the fundamental level of semantic embeddings. We use the exact wording from the PVQ-RR questionnaire specifications (see Subsection A.7) for those models that allow including a prompt [^2] We run all models on CPU to embed all 57 items in less than 20 minutes on a standard ThinkPad T480 from 2018.. The PVQ-RR is available in gendered version using different pronouns. We account for this feature by creating an average of the female and male item embeddings in order to reduce possible gender biases.
@@ 4.2 Survey and Questionnaire Item Embeddings Differentials (SQuID)
Working with raw embeddings as directly retrieved from the models, we faced the same problem of missing negative correlations as the other researchers whose work we discussed in Section 2 (see the similarity matrix on the left of Figure 7). As we will use MDS for dimensionality reduction which crucially builds on preserving distances, it was especially important for us to find a way to make the signal of negative correlations stronger.
As discussed before, the root of the problem lies in the high-dimensionality of item embeddings that encode many common linguistic features of natural language use which artificially boost similarity between items. To create a vector that captures this unneeded information, we take an average embedding over all items (a "questionnaire embedding") and subtract it from each of the item embeddings: $\mathbf{y}_{i}-\overline{\mathbf{y}}$ for item $i$, where $\overline{\mathbf{y}}$ is an average over all 57 items of the PVQ-RR: $\overline{\mathbf{y}}=\frac{1}{57}\sum_{i=1}^{57}\mathbf{y}_{i}$. This straightforward method is inspired by the early examples of vector arithmetic demonstrated with Word2Vec Mikolov et al. (2013) and the more recent idea of community embeddings Waller and Anderson (2021). As can be seen in the right hand part of Figure 7, this approach is simple yet very effective in making the signal of negative correlations appear without any domain-specific finetuning or other adaption procedures. SQuID works with any model and embeddings of arbitrary dimensions, offering a high degree of flexibility. We believe that a substantial part of the under-performance of embedding approaches compared to human data in other studies, is due to methodological reasons and can be overcome with SQuID.
@@
Table 1: Internal consistency as measured through Cronbach’s alpha for each of the 19 value dimensions and the average over dimensions. Rowname abbreviations refer to the models covered in Subsection 4.1 (all embeddings were SQuID treated). We report a random baseline of embeddings created from a normal distribution with mean zero and unit variance (averages over 10 000 samples). Human data is retrieved from Table S2 in SI of Schwartz and Cieciuch (2022). Linq-Embed-Mistral is generally the best performing embedding model with an average value for Cronbach’s alpha that is above even the human benchmark. Values close to 0.8 or above for most value dimensions point to adequate internal consistency (with notable exceptions such as HUM/humility, which is also low in the human data). For fully spelled out dimension names in the column header see Figure 7.